OpenStudy (anonymous):

find the amplitude,period, horizontal shift and vertical shift of the following: y=2cos(2pix) and y=sin(x-pi/4)+2

jimthompson5910 (jim_thompson5910):

how far did you get?

OpenStudy (anonymous):

well for the first one i got that the amplitude is 2, the period is 1, the horizontal shift to be none and the vertical shift to be none. For the second one i did not get the amplitude or period but for the horizontal shift i got pi/4 units to the right and the vertical shift to be 2 units up

jimthompson5910 (jim_thompson5910):

very good so far

jimthompson5910 (jim_thompson5910):

the only thing you're missing is the period of y=sin(x-pi/4)+2

jimthompson5910 (jim_thompson5910):

y=sin(x-pi/4)+2 is the same as y=sin(1x-pi/4)+2

jimthompson5910 (jim_thompson5910):

The coefficient of the x term is 1, so B = 1 Period: T = 2pi/B T = 2pi/1 T = 2pi which means the period of y=sin(x-pi/4)+2 is 2pi

jimthompson5910 (jim_thompson5910):

Side Note: I'm using the basic form y = A*sin(Bx - C) + D

OpenStudy (anonymous):

ok for the amplitude of the second function is it one?

jimthompson5910 (jim_thompson5910):

That is correct. A = 1 So Amplitude = |A| = |1| = 1

OpenStudy (anonymous):

thank you soo much! i have to go now unfort. i have to go to the store but thanks for the help i might be on later if you are on would you mind helping me more?

jimthompson5910 (jim_thompson5910):

sure I'd love to

OpenStudy (anonymous):

ok thanks message you after a bit

jimthompson5910 (jim_thompson5910):

alright, cya later